# NRSurNN3dq4: A Deep Learning Powered Numerical Relativity Surrogate for Binary Black Hole Waveforms

### v1

@@ III.2 Network architecture and training
Conceptually, we propose to build a neural network that maps the 3 physical input parameters (mass ratio and dimensionless individual spins) to the 100-dimensional orthogonal space corresponding to the chosen PCA bases. We can then apply the inverse PCA transform and obtain the reconstructed waveform $h$. With this formulation, we can now define our optimization objective, or loss function. On the one hand, we want to ensure the accuracy of the predicted PCA components. This can be quantified using a mean absolute error (MAE) criterion,
@@
where $c$ is the generated array of the 100 PCA components for each generated waveforms, and $\hat{c}$ represents the corresponding ground truth values of these components. We choose to use MAE rather than its square as it is more robust to outliers. On the other hand, from a practical perspective, we eventually want to maximize the overlap $\mathcal{O}$ between the waveforms themselves, as that is the usual evaluation metric for approximants. It should be useful, then, to make sure our loss function includes information about the overlap. Note that since we want to maximize the overlap, we should then minimize the mismatch, given by
@@
is the inner product between waveforms. However, it is more convenient to work with the base-10 logarithm of this quantity, as this will keep the magnitude of the loss (and gradients) more consistent as the mismatch reaches lower values. As such, our loss function $\mathcal{L}$ shall be
@@
Figure 4: (a) Training and validation loss on the approximant dataset and, (b) validation mismatches on the approximant dataset.
The network used is a simple multi-layered perceptron with 3 hidden layers, with 64, 512 and 1024 neurons, respectively. The model architecture is illustrated in Figure 3. In this work, all computations make use of the NVIDIA V100 32GB GPU nodes provided by the Artemisa cluster at the University of València. The number and width of the hidden layers are the result of an optimization process through the optuna package optuna_2019 (note that in the case of the layer width, we optimized over powers of $2$, in order to reduce redundancy). The model is then pre-trained for 20,000 epochs in the approximant dataset, with an 80/20 train/validation split. We use the Sophia optimizer sophiaG , with an initial learning rate of $3\times 10^{-3}$ (obtained through the learning rate finder process outlined in lrfinder ), and a learning rate scheduler reducing the learning rate by a factor of 4 after 1000 epochs without improved performance on the validation set. During the training loop we monitor the mean value of the validation mismatches, and save the model every time there is an improvement (that is, when the mean validation mismatch decreases in comparison to the previous best value), guaranteeing we have the best performing model on the training set after all epochs are done. This model will then be fine-tuned on the numerical relativity data. Since the numerical relativity dataset is quite small, we proceed as follows: first, we set apart 1/8th of the waveforms for testing purposes. With the remaining waveforms, we use k-fold cross validation by splitting the data in 5 equal-sized subsets of “folds”. For each iteration, we consider the non-chosen folds as a training set, and the chosen fold as the validation. This allows us to have 5 iterations of training, using the model trained in the approximant as a starting point, and using a lower learning rate than in the pre-training step ($1\times 10^{-4}$) in order to avoid an excessive deviation from the region of the space of network weights arrived at in pre-training. When this process is completed, the weights of the 5 trained models are averaged to obtain a final model, which is then evaluated on the test set. This process should be seen as a sort of weight averaging, which has been shown to lead to improved generalization performance weight_avg .
Figure 5: Distribution of validation mismatches for the model trained on the approximant dataset.
@@ IV Results
Fold 0 Fold 1 Fold 2 Fold 3 Fold 4 Avg. Weights No fine-tuning
Mean mismatch $5.57{\times}10^{-3}$ $6.50{\times}10^{-3}$ $5.72{\times}10^{-3}$ $5.81{\times}10^{-3}$ $7.15{\times}10^{-3}$ $\mathbf{5.43{\times}10^{-3}}$ $2.09{\times}10^{-2}$
Max mismatch $2.57{\times}10^{-2}$ $4.39{\times}10^{-2}$ $2.16{\times}10^{-2}$ $2.92{\times}10^{-2}$ $3.16{\times}10^{-2}$ $\mathbf{2.10{\times}10^{-2}}$ $1.32{\times}10^{-1}$
Table 1: Mean and maximum mismatch values for the various models in the set-aside test set
@@ IV.1.1 Results on the approximant waveforms
Figure 4(a) shows the evolution of the loss function on the training and validation sets as the model trains. We can see that the training loss has some moments of rapid drops, but otherwise decreases steadily. The validation loss follows this same pattern for about the first 7500 epochs, but then starts to plateau, showing very minor improvements until the end. Correspondingly Fig. 4(b) shows the evolution of the mismatches on the validation set. The lowest mean validation mismatch is achieved at epoch 19910, and as such the final saved model state corresponds to this epoch.
In Fig. 5 we depict the distribution of the mismatch metric on the validation set after pretraining on the approximant dataset. The distribution is peaked around $10^{-5}$, with the worst mismatch found in the validation set being $9.5\times 10^{-3}$. This indicates that our network does a good job as a surrogate model of the NRSur7dq4 approximant under the parameter restrictions of our setup.
@@ IV.1.2 Results on the Numerical Relativity waveforms
Figure 6: Distribution of mismatches on the set-aside test dataset, shown for the fine-tuned and weight-averaged final model, as well as for the model trained only on the approximant.
Table 1 shows the mean and maximum mismatch values for the various trained models in the set-aside test set, including a model trained only on the approximant, before fine-tuning. We can see that the weight averaging of the 5 models trained on the folds results in better mismatches compared to models trained on each individual fold. We can also see that skipping the fine-tuning step (last column) results in mismatches an order of magnitude larger than the fine-tuned models, as further made evident in Figure 6. This makes it clear that the fine-tuning step is crucial to obtain a good numerical relativity surrogate with our methodology.
It is also necessary to point out that the results on the numerical relativity data are worse than the results on the approximant data alone. This may be explained by the increased complexity of the numerical relativity waveforms, as well as by the relatively small amount of data available. Nonetheless, the model is still accurate enough for parameter estimation, as we shall show in Section IV.3.
@@ IV.2 Speed
Figure 7: Cumulative waveform generation time for NRSurNN3dq4.
@@
To test the runtime performance of the model, we generate $1{\times}10^{7}$ randomly sampled waveforms in batches of $1{\times}10^{5}$ using an NVIDIA V100 GPU. Note that in this process we are now including the scaling from geometric to physical units. Figure 7 shows the cumulative time taken by NRSurNN3dq4 to generate the waveforms. We must note here that there is a warm-up period, typically around 500 milliseconds of overhead, attributable to CUDA initialization. However, once initialized, the total generation time for $10^{7}$ waveforms is achieved in under 40 milliseconds.
Figure 8 presents a box plot of the generation time for sets of 20 batches of size $10^{2}$ to $10^{5}$. As may be expected, as long as GPU memory is not exceeded, the batch size does not significantly affect the mean generation time. This means that approximants in the vein of NRSurNN3dq4 can be used to generate massive GW template banks in very short time frames.
@@ IV.3 Parameter estimation
As a way of exploring the ability of NRSurNN3dq4 to be used in parameter estimation pipelines, we here discuss a basic test we conducted using the bilby software suite for GWs bilby_paper . We must note that the current sampler and likelihood implementations in bilby do not allow us at this point to take advantage of CUDA parallelization of the likelihood (though it is in principle possible to write a custom sampler that would manage this). Therefore, we will be running the model in the CPU for this section, focusing mostly on the accuracy of the model for parameter estimation. For comparison’s sake, running the model in CPU mode, it takes us 103 ms to generate a 1000 waveforms in serial. On the other hand, in the same machine, it takes us 810 ms seconds to generate 1000 waveforms with NRSur7dq4.
For this experiment, we will focus on performing a parameter inference run on the GW200311_115853 event from the third Gravitational-Wave Transient Catalog GWTC-3 , using the nessai sampler nessai and NRSurNN3dq4 as the waveform model. This model was chosen as an illustrative example due to its relatively high SNR, but performance in other events can be expected to be similar. With 300 live points, this run took 6 minutes and 2 seconds to complete, with 390,704 likelihood evaluations accounting for 39.5 seconds of that time. In Fig. 9 we show the results of the PE run, with the posteriors being calculated for the mass ratio $q$ and the spins $\chi_{1},~{}\chi_{2}$, as well as the chirp mass $\mathcal{M}=\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}$, luminosity distance $d_{L}$, the sky position angular coordinates (right ascension $\alpha$ and declination $\delta$), and the polarization angle $\Psi$. All estimated values are contained in the 90% credible regions of the LVK posterior distributions, showing that our model can be used to reliably reconstruct physical parameters in line with LVK standards.
@@ V Conclusions
In this paper we have introduced NRSurNN3dq4, a neural-network-based surrogate model designed to rapidly and accurately generate non-precessing BBH waveforms. By mapping the physical input parameters to a reduced 100-dimensional space of PCA coefficients, our model can reconstruct waveforms with low mismatches. The initial results on approximant data show average waveform mismatches of the order of $10^{-5}$, yielding immediate promise for these methods as reduced-order models for approximants. While the results after fine-tuning on numerical relativity data are not quite as good as for the case of approximants, likely due to the lack of data, the average numerical relativity waveform mismatches are of the order of $5{\times}10^{-3}$. The fact that the fine-tuned model achieves mismatches with numerical relativity waveforms an order of magnitude lower than the pretrained model reinforces the idea that this 2-step training procedure is effective. Crucially, timing tests demonstrate that NRSurNN3dq4 can generate over a million waveforms in under a tenth of a second when implemented in a GPU. This makes NRSurNN3dq4 a powerful tool when the generation of large amounts of waveforms in parallel is required, such as building template banks, building datasets for training machine learning models, or for GPU-parallelized parameter estimation.
The evaluation of this model within the bilby framework shows results that are consistent with those reported in the literature. The results for parameter estimation confirm the model’s accuracy and open the door for further integration. A GPU adaptation of existing bilby samplers and likelihood implementations in order to leverage the ability to generate arbitrarily large numbers of waveforms has the potential to speed up parameter estimation in a very significant way and must be the focus of future work. Additional future efforts could focus on extending the model’s capabilities to precessing systems and higher modes, as well as refining its accuracy through the exploration of alternative dimensionality reduction methods and alternative network architectures.

### v3

@@ III.2 Network architecture and training
Conceptually, we propose to build a neural network that maps the 3 physical input parameters (mass ratio and dimensionless individual spins) to the 42-dimensional orthogonal space corresponding to the chosen PCA bases. We can then apply the inverse PCA transform and obtain the reconstructed waveform $h$. With this formulation, we can now define our optimization objective, or loss function. On the one hand, we want to ensure the accuracy of the predicted PCA components. This can be quantified using a mean absolute error (MAE) criterion,
@@
where $c$ is the generated array of the N PCA components (N=42 in our case) for each generated waveforms, and $\hat{c}$ represents the corresponding ground truth values of these components. We choose to use MAE rather than its square as it is more robust to outliers. On the other hand, from a practical perspective, we eventually want to maximize the overlap $\mathcal{O}$ between the waveforms themselves, as that is the usual evaluation metric for approximants. It should be useful, then, to make sure our loss function includes information about the overlap. Note that since we want to maximize the overlap, we should then minimize the mismatch $\mathfrak{M}$, given by
@@
is the inner product between waveforms (note that if this operation is being applied to compare a waveform with detector data instead, the integral limits should change to the relevant frequency range). While the mismatch is often optimized over time and phase, the alignment of the waveforms and the setting of the initial phase in our datasets make this unnecessary. Numerically, it will be more convenient to work with the base-10 logarithm of this quantity, as this will keep the magnitude of the loss (and gradients) more consistent as the mismatch reaches lower values. As such, our loss function $\mathcal{L}$ shall be
@@
Figure 4: Validation loss on the approximant datasets during pre-training.
The network used is a simple multi-layered perceptron with 3 hidden layers, with 64, 512 and 1024 neurons respectively, using GELU activation functions. The model architecture is illustrated in Figure 3. In this work, all computations make use of the NVIDIA H100 GPU nodes provided by the Artemisa cluster at the University of València. The number and width of the hidden layers are the result of an optimization process through the optuna package [53] (note that in the case of the layer width, we optimized over powers of 222, in order to reduce redundancy). The model is then pre-trained for up to 10,000 epochs for each approximant dataset, with an 80/20 train/validation split. We use the Sophia optimizer [54], with an initial learning rate obtained through the learning rate finder process outlined in [55], and a learning rate scheduler reducing the learning rate by a factor of 2 after 500 epochs without improved performance on the validation set. During the training loop we monitor the mean value of the validation mismatches, and save the model every time there is an improvement (that is, when the mean validation mismatch decreases in comparison to the previous best value), guaranteeing we have the best performing model on the training set after all epochs are done. This model will then be fine-tuned on the numerical relativity data. Since the numerical relativity dataset is quite small, we proceed as follows: first, we set apart 1/8th of the waveforms for testing purposes. With the remaining waveforms, we use k-fold cross validation by splitting the data in 5 equal-sized subsets or “folds”. For each iteration, we consider the non-chosen folds as a training set, and the chosen fold as the validation. This allows us to have 5 iterations of training, using the model trained in the approximant as a starting point, and using a lower learning rate than in the pre-training step ($1\times 10^{-4}$) in order to avoid an excessive deviation from the region of the space of network weights arrived at in pre-training. When this process is completed, the weights of the 5 trained models are averaged to obtain a final model, which is then evaluated on the test set. This process should be seen as a sort of weight averaging, which has been shown to lead to improved generalization performance [56].
Figure 5: Distribution of parameters in the NR dataset’s train/validation and testing sets.
Figure 6: Distribution of validation mismatches for the models trained on the approximant datasets.
Figure 7: Heatmaps showing the mean validation mismatch of the pre-trained network per bin, as a function of the (inverted) mass ratio and the effective inspiral spin. Top left: NRHybSur3dq8. Top right: SEOBNRv5HM. Bottom left: SEOBNRv4HM_PA. Bottom right: IMRPhenomTHM.
@@ IV.1.1 Results on the approximant waveforms
In Figure 6 we depict the mismatch distributions for each of the four individual waveform approximants. The NRHybSur3dq8 model yields a broad distribution, featuring a significant population with mismatches below $10^{-6}$ but also a flat feature near $10^{-4}$, with a tail extending to $2.5{\times}10^{-2}$. The distribution for SEOBNRv5HM peaks around $10^{-5}$, with its worst mismatches reaching $3.4{\times}10^{-3}$. Lastly, the SEOBNRv4HM_PA and IMRPhenomTHM models display very similar and more concentrated distributions, both showing a ramp up to the peak around $10^{-3}$, with a sharp drop off to maximum mismatches of order $10^{-2}$. Figure 7 complements this information by showing the mean mismatch value per bin over the plane defined by the inverted mass ratio and the effective inspiral spin $\chi_{\rm eff}=\frac{\chi_{1}+q\chi_{2}}{1+q}$ for each of the approximants. The similarity between SEOBNRv4HM_PA and IMRPhenomTHM is once again made evident, with both showing mostly uniform mean values between $10^{-4}$ and $10^{-3}$. Given that both these approximants are used through their LALSimulation implementation, it is possible that there is some common processing in the generation of these waveforms. The network trained on the SEOBNRv5HM approximant data also shows fairly uniform behaviour across the parameter space, with some higher mismatches appearing at higher $\chi_{\rm eff}$ values, but otherwise remaining between $10^{-6}$ and $10^{-4}$. The behaviour on the NRHybSur3dq8 data stands out as most interesting, with the appearance of some strong heterogeneity over the parameter space, with particular expression near the origin, where a diagonal line from $(\frac{1}{q}=0.25,\chi_{\rm eff}=-0.7)$ to $(\frac{1}{q}=0.125,\chi_{\rm eff}=0.3$) shows mean mismatches up to three orders of magnitude larger than its surrounding regions. It is unclear why this happens and visual inspection of the generated waveforms does not show any collapse behavior. Nonetheless, even in these cases the mismatches are decent and we do not expect it to affect the fine-tuning process significantly.
@@ IV.1.2 Results on the Numerical Relativity waveforms
Figure 8: Distribution of mismatches on the set-aside test NR dataset, shown for the fine-tuned and weight-averaged models, as well as for the pre-trained-only models, along with the mismatches of the base approximants themselves with NR.
Table 1 shows the mean and maximum mismatch values for the various trained models in the set-aside NR test set. We can see that the weight averaging of the 3 models trained on the folds seems to achieve the goal of enforcing generalization capacity. While not always the top performing model, the weight-averaged models always avoid having large maximal mismatches, while retaining mean performance. This lines up with the theoretical advantages of weight-averaging regarding model generalization. In Figure 8 we show the mismatches for the fine-tuned models’ reconstructions of NR test set waveforms, when pre-trained on different datasets. We also show the mismatches of the corresponding model after the pre-training step, as well as the mismatches obtained using the base approximant used to train the model. We should start by pointing out the difference in accuracy to NR across the different approximants, with the NRHybSur3dq8 approximant achieving the best median mismatches, around $4{\times}10^{-5}$, and the SEOBNRv4HM_PA approximant being the relatively worst performer with median mismatches of $5.3{\times}10^{-3}$, two orders of magnitude larger than the best performer. In all cases, the median mismatches of the pre-trained models correspond remarkably well with the mismatches of the respective base approximants, showing that the proposed architecture can accurately reproduce the outputs of GW approximants. The fine-tuning process yields very different results depending on the base approximant used for pretraining. In the NRHybSur3dq8 case, the major gain with the fine-tuning process is the suppression of a tail that previously extended beyond $10^{-1}$, arriving at a final distribution with a median of $2.2{\times}10^{-5}$, smaller than the median from the approximant by almost a factor of 2. For the other approximants, however, the fine-tuning process shows gains in the median mismatch between one and two orders of magnitude, with every fine-tuned model achieving median mismatches below $8{\times}10^{-5}$, matching the accuracy from the base NRHybSur3dq8 approximant, and in the case of the model pre-trained on IMRPhenomTHM, surpassing it slightly. The convergence to this range of values regardless of the accuracy of the base approximant suggests that the strategy used to train DANSur3dq8  can successfully use synthetic data obtained from approximations as a springboard for learning to reproduce simulation data, with accuracy gains that can span multiple orders of magnitude over the initial approximation.
Model Config Mean Max
NRHybSur3dq8 W. Avg. $2.3\times 10^{-5}$ $2.7\times 10^{-3}$
Fold 1 $2.8\times 10^{-5}$ $\mathbf{2.0\times 10^{-3}}$
Fold 2 $\mathbf{2.0\times 10^{-5}}$ $1.0\times 10^{-2}$
Fold 3 $3.3\times 10^{-5}$ $3.0\times 10^{-3}$
SEOBNRv5HM W. Avg. $5.3\times 10^{-5}$ $\mathbf{2.8\times 10^{-3}}$
Fold 1 $\mathbf{4.7\times 10^{-5}}$ $1.0\times 10^{-2}$
Fold 2 $5.8\times 10^{-5}$ $5.8\times 10^{-3}$
Fold 3 $8.3\times 10^{-5}$ $5.7\times 10^{-3}$
SEOBNRv4HM_PA W. Avg. $6.2\times 10^{-5}$ $\mathbf{6.4\times 10^{-3}}$
Fold 1 $\mathbf{4.9\times 10^{-5}}$ $4.0\times 10^{-2}$
Fold 2 $7.5\times 10^{-5}$ $1.3\times 10^{-2}$
Fold 3 $6.0\times 10^{-5}$ $6.6\times 10^{-3}$
IMRPhenomTHM W. Avg. $\mathbf{2.7\times 10^{-5}}$ $4.9\times 10^{-3}$
Fold 1 $3.5\times 10^{-5}$ $1.4\times 10^{-2}$
Fold 2 $3.8\times 10^{-5}$ $2.2\times 10^{-2}$
Fold 3 $3.2\times 10^{-5}$ $\mathbf{2.7\times 10^{-3}}$
Table 1: Mean and maximum mismatch values for the various models in the set-aside test set, showing results for individual folds and the weight-averaged model
@@ IV.2 Speed
Figure 9: Cumulative waveform generation time for DANSur3dq8.
@@
To test the runtime performance of DANSur3dq8  models, we generate $1{\times}10^{7}$ randomly sampled waveforms in batches of $1{\times}10^{5}$ using an NVIDIA H100 GPU. Note that in this process we are now including the scaling from geometric to physical units. Figure 9 shows the cumulative time taken by DANSur3dq8 to generate the waveforms. We must note here that there is a warm-up period, typically around 500 milliseconds of overhead, attributable to CUDA initialization. However, once initialized, the total generation time for $10^{7}$ waveforms is achieved in under 40 milliseconds.
Figure 10 presents a box plot of the generation time for sets of 20 batches of size $10^{2}$ to $10^{5}$. As may be expected, as long as GPU memory is not exceeded, the batch size does not significantly affect the mean generation time. This means that approximants in the vein of DANSur3dq8 can be used to generate massive GW template banks in very short time frames.
@@ IV.3 Parameter estimation
As a way of exploring the ability of DANSur3dq8 to be used in parameter estimation pipelines, we here discuss a basic test we conducted using the bilby software suite for GWs [15]. We must note that the current sampler and likelihood implementations in bilby do not allow us at this point to take advantage of CUDA parallelization of the likelihood (though it is in principle possible to write a custom sampler that would manage this). Therefore, we will be running the model in the CPU for this section, focusing mostly on the accuracy of the model for parameter estimation. To this end, we make use of pytorch (v2.4.1) model serialization through TorchScript, allowing for better performance and ease of use. For comparison’s sake, running the model on CPU, implemented through the gwsurrogate api, it takes us 0.3 s on average (over 10 runs) to generate 1000 waveforms in serial. On the other hand, in the same machine, it takes us 5.6 s seconds on average to generate 1000 waveforms with the  NRHybSur3dq8 implementation in gwsurrogate (v1.1.6). We have restricted the generation with NRHybSur3dq8  to the (2,2) mode, and kept the same time grid as DANSur3dq8. These timing tests have been run on an AMD EPYC 9454 48-core CPU at the Artemisa cluster of the University of Valencia. For this experiment, we will focus on performing a parameter inference run on the GW200311_115853 event from the third Gravitational-Wave Transient Catalog [14], using the nessai sampler [22] and the DANSur3dq8  pretrained on the NRHybSur3dq8 approximant as the waveform model. This event was chosen as an illustrative example due to its relatively high SNR, but performance in other events can be expected to be similar. With 1000 live points, this run took 3 minutes and 35 seconds to complete, with 346,596 likelihood evaluations accounting for 15.8 seconds of that time. In Fig. 11 we show the results of the PE run, with the posteriors being calculated for the mass ratio $q$ and the spins $\chi_{1},~\chi_{2}$, as well as the chirp mass $\mathcal{M}=\frac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}$, luminosity distance $d_{L}$, the sky position angular coordinates (right ascension $\alpha$ and declination $\delta$), and the polarization angle $\Psi$. All estimated values are contained in the 90% credible regions of the LVK posterior distributions, showing that our model can be used to reliably reconstruct physical parameters in line with LVK standards. These results do have the caveat that our model does not include higher harmonics, which are more important the larger the mass ratio is, as well as allowing for better precision at higher SNR values. The lack of precession modeling must also be taken into account. This will bias our model under certain conditions.
@@ V Conclusions
In this paper we have introduced a 2-step training strategy to generate neural-network-based NR surrogate models (DANSur3dq8), shown to rapidly and accurately generate non-precessing BBH waveforms.
In a first step, we use BBH approximants to generate large datasets of gravitational waveforms for four different approximants, using the resulting datasets to pre-train surrogate networks. We show that, after the pretraining step, the resulting model can already accurately reproduce the outputs of the base approximant. Together with our usage of a reduced-order PCA basis, this step effectively restricts the vast function space of the neural network, so that when we start our second step, our initial point in the network’s state space is one which produces gravitational waveforms from BBH systems. The second step itself consists of fine-tuning a pre-trained network on waveforms from NR simulations, allowing the training process to adjust the basis through backpropagation as well. This results in median mismatches around $5{\times}10^{-5}$ on average, regardless of the approximant used for pretraining. When compared to the base approximant used for pre-training, we show that the fine-tuned DANSur3dq8 model can improve mismatches with NR waveforms by up to two orders of magnitude, reinforcing the effectiveness of the 2-step training procedure.
Timing tests demonstrate that surrogates trained under this strategy can generate over a million waveforms in under a tenth of a second when implemented in a GPU. This makes them a powerful tool when the generation of large amounts of waveforms in parallel is required, such as building template banks, building datasets for training machine learning models, or for GPU-parallelized parameter estimation.
Despite the studied case’s limitations, namely in the absence of higher modes and precession, as well as the limited number of cycles used, the evaluation of the DANSur3dq8 model within the bilby framework shows results that are consistent with those reported in the literature. The results for parameter estimation confirm the model’s accuracy and open the door for further integration. A GPU adaptation of existing bilby samplers and likelihood implementations in order to leverage the ability to generate arbitrarily large numbers of waveforms has the potential to speed up parameter estimation in a very significant way and must be the focus of future work. Additional future efforts will focus on augmenting DANSur3dq8’s capabilities to model precessing systems and higher modes, refining its accuracy through the exploration of alternative dimensionality reduction methods and alternative network architectures, and extending the parameter space.